OpenStudy (lilai3):

Please help! It will be greatly appreciated! Will award medal! Mrs. Valdez can assemble a chair in 1 day and a table in 4 days. Graph the function 1x+4y=20 to determine how many of each type of furniture Mrs. Valdez can assemble in 20 days.

OpenStudy (clalgee):

So, 1 chair- 1 day and 1 table= days right?

OpenStudy (clalgee):

That will be 25 pieces or furniture 20 days. How? 20 Chairs= 20 Days and 5 Tables= 4 days. So 20 chairs + 5 tables= 25 furniture in total

OpenStudy (lilai3):

well, can you graph it though..or try to? well, at least make the chart (the function chart) please.

OpenStudy (clalgee):

ok, give me one second

OpenStudy (lilai3):

okay

OpenStudy (clalgee):

I did the first part. Just finish the next couple

OpenStudy (clalgee):

double click on the chart to edit it

OpenStudy (lilai3):

oh gee. thanks.uhm. so the slope must be 1/4... right?

OpenStudy (clalgee):

correct

OpenStudy (lilai3):

oh okay thank you for your help i get it now. thank you so much.

OpenStudy (lilai3):

would you be willing to help me on another problem?

OpenStudy (clalgee):

No problem and sure. What's your question?

OpenStudy (lilai3):

A fitness center has set a goal to have 500 members. The fitness center already has 150 members and adds an average of 25 members per month. The function f(x)=25x +150 represents the membership after x months. Graph the function to determine the number of months it will take for the fitness center to reach its membership goal.

OpenStudy (clalgee):

Well, here is the equation: 1x + 4y = 20 4y = 20 - 1x y = (20 / 4) - 1/ 4x so the equation for a graph is: y= 5 - 1 / 4x As well as graphing it, hold on right there.

OpenStudy (clalgee):

It will look something like this

OpenStudy (lilai3):

i'm sorry but like i still don't get it